# Hardware-Aware Parallel Prompt Decoding for Memory-Efficient Acceleration of LLM Inference

### v2

@@ Abstract
The auto-regressive decoding of Large Language Models (LLMs) results in significant overheads in their hardware performance. While recent research has investigated various speculative decoding techniques for multi-token generation, these efforts have primarily focused on improving processing speed such as throughput. Crucially, they often neglect other metrics essential for real-life deployments, such as memory consumption and training cost. To overcome these limitations, we propose a novel parallel prompt decoding that requires only % trainable parameters, enabling efficient training on a single A100-40GB GPU in just 16 hours. Inspired by the human natural language generation process, PPD approximates outputs generated at future timesteps in parallel by using multiple prompt tokens. This approach partially recovers the missing conditional dependency information necessary for multi-token generation, resulting in up to a 28% higher acceptance rate for long-range predictions. Furthermore, we present a hardware-aware dynamic sparse tree technique that adaptively optimizes this decoding scheme to fully leverage the computational capacities on different GPUs. Through extensive experiments across LLMs ranging from MobileLlama to Vicuna-13B on a wide range of benchmarks, our approach demonstrates up to 2.49 speedup and maintains a minimal runtime memory overhead of just . More importantly, our parallel prompt decoding can serve as an orthogonal optimization for synergistic integration with existing speculative decoding, showing up to further speed improvement. Our code is available at https://github.com/hmarkc/parallel-prompt-decoding.
@@ 1 Introduction
The recent advances in large language models (LLMs) are increasingly gaining influence across various AI applications. However, autoregressive generation, the de facto approach employed in LLM inference, suffers from inadequate hardware performance due to its inherent sequential nature [23]. Speculative decoding [13, 2, 11], an emerging acceleration technique, employs a guess-and-verify framework for LLM inference, where a smaller draft model first predicts multiple tokens sequentially and then the original LLM verifies them in parallel. Despite its potential, the effectiveness of speculative decoding is limited by the complexity and cost of training a draft model capable of consistently achieving high acceptance rates across diverse base models and datasets. Additionally, the extra runtime memory overhead for executing draft models poses a significant barrier to the broader adoption of speculative decoding, particularly in edge and mobile environments where memory capacity is limited. Considering the growing need for user privacy and personalization, deploying LLMs on devices urges a more memory- and cost-efficient solution for accelerating LLM inference. Recent efforts have explored the possibility of generating multiple tokens in parallel without relying on a separate transformer draft model [20]. Approaches such as inserting additional decoding heads [1] and retrieving frequently used tokens [9] are employed to enhance performance. However, these methods either aggressively assume conditional independence among the tokens generated in a single step [1, 9], or use placeholder tokens (e.g., [PAD] token) that do not convey enough contextual information [20]. Therefore, they often suffer from low acceptance rates or degradation in output quality due to the lack of sufficient conditional information during inference.
@@
To alleviate the complexity and overhead associated with the use of draft models while maintaining a high acceptance rate, we propose Parallel Prompt Decoding (PPD), a novel architecture-agnostic and memory-efficient framework that adopts prompt tuning for non-autoregressive LLM inference. Inspired by the human natural language generation process where continuous words like common expressions and phrases are produced simultaneously, PPD introduces the use of prompt tokens, the meticulously trained embeddings, for multi-token prediction. Specifically, these trained prompt tokens are appended to the original input sequence in parallel, enabling the concurrent generation of multiple output tokens in a single forward pass. The key intuition of PPD lies in the observation that if trained properly, prompt tokens appended to the input can approximate tokens generated at future timesteps, thereby partially recovering the missing conditional dependency information for multi-token generation. By strategically positioning trained prompt tokens, PPD achieves up to a 28% higher acceptance rate when predicting long-range tokens. To further increase the token acceptance rate, we generate multiple candidate continuations with each prompt token and use them in combination with a customized tree attention mask to minimize the computation and memory overhead. The capability of PPD to use low-cost prompt tokens for accurate multi-token prediction forms the foundation for accelerating LLM inference. As shown in Figure 1, PPD achieves a comparable speedup to the state-of-the-art speculative decoding approaches with negligible memory overhead and reduced training cost. Moreover, to facilitate the optimized implementation of PPD across different hardware platforms, we propose a hardware-aware dynamic sparse tree technique that adaptively refines the prompt structure during runtime based on the computational resources available on the specific hardware.
@@
To demonstrate the effectiveness of our approach, we evaluate PPD on MobileLLaMA [6], Vicuna-7b and Vicuna-13b [5]. Running on a single GPU using the A100-40GB and RTX 4090, our method achieves a speedup ratio for inference from 2.12 to 2.49 across a diverse range of popular datasets including MT-Bench, HumanEval, and GSM8K. Our experiments demonstrate that PPD not only achieves comparable throughput to the state-of-the-art speculative decoding method, but it also manages this with significantly fewer trainable parameters—specifically, 0.0002% of trainable parameters—and incurs only a minimal memory overhead (0.0004%), showcasing that PPD is remarkably cost- and memory-efficient. The training of prompt tokens can be completed in 16 hours using one A100 GPU, 8 hours using four GeForce RTX 3090 GPUs, compared to the 1-2 days on four A100 GPUs required for Eagle [16]. Furthermore, since PPD does not require the modification of the original LLM or the addition of extra networks, it is highly adaptable and orthogonal to other decoding techniques. For instance, it can be effectively combined with a draft model to further reduce inference latency.
@@
A hardware-aware dynamic sparse tree technique that adaptively optimizes the prompt structure of PPD at runtime based on the available compute and memory resources, facilitating its efficient deployment on various hardware platforms.
@@ 2.1 Speculative Decoding
The guessing phase of speculative decoding adopts a lightweight draft model to generate multiple tokens at an increased speed [11]. During the verification stage, the original LLM subsequently determines the acceptance of the guessed tokens. It is worth noting that both draft and original models still follow the auto-regressive inference scheme. The speedup comes from two factors: i) the draft model runs much faster than the original model and more tokens can be generated within the same time unit; and ii) token verification is executed concurrently, either by batching or by incorporating multiple candidates into a single input using customized sparse attention masks [18]. Therefore, the overall speedup depends on the acceptance rate and the inference latency of draft models.
Building on the speculative decoding scheme, various studies have been conducted to further optimize its inference speed. To improve the accuracy of the draft model and its token acceptance rate, Eagle [16] incorporates the hidden features into the draft model’s forward pass. SpecInfer [18] adopts a tree-based speculative inference and verification scheme, improving the diversity of speculation candidates. Sequoia [4] optimizes the sparse tree structure by considering the capability of the underlying hardware platforms. However, most of these methods require the storage and maintenance of a separate draft model. Moreover, there is extra complexity in designing an efficient draft model.
@@ 2.2 Parallel Decoding
To overcome the inherent limitations of autoregressive inference and the memory overhead associated with using a separate draft model, several attempts have been made to integrate both guessing and verification using one unified model. Medusa111We categorize Medusa as parallel decoding because it only adopts LM heads instead of separate models. [1] introduces language model (LM) heads at the final layer of the original LLM, facilitating the generation of multiple tokens in a single forward pass. It also utilizes tree attention masks in its verification process to increase speed even further. To enhance token drafting with retrieval-augmented generation [10], Rest [9] introduce retrieval-based decoding tailored for specific scenarios. Inspired by Jacobi decoding [20] that adopts multiple special tokens to accelerate machine translation, Lookahead Decoding [8] improves upon this method by generating parallel n-grams and employing a caching memory pool. To capture more information while using multiple special tokens at distinct positions, PaSS [19] trains additional tokens with embedding layers for parallel decoding. Hierarchical parallel decoding [17] introduces the use of and tokens, enabling parallel execution of multiple structural subroutines.
Our approach can be categorized as parallel decoding, with three novel features to distinguish it from other approaches: 1) PPD trains the embeddings of parameterized ensemble prompt tokens, 2) it utilizes a dynamic sparse tree, adapting its structure at every inference step, and 3) we propose a hardware-aware algorithm for designing a dynamic sparse tree tailored to each hardware platform.
@@ 3 Parallel Prompt Decoding (PPD)
PPD trains embeddings for prompt tokens rather than developing a separate model. Our method integrates three substeps into a single decoding step, following the guess-and-verify strategy: (1) candidate generation, where multiple candidate continuations222A candidate token, also referred to as a ”guess token”, is a draft token generated from a prompt token. are predicted by strategically inserting the prompt tokens into the input sequence. Tree attention [18] merges the processing of multiple candidates into a single forward pass; (2) candidate verification, where two verification schemes, exact matching [8] and typical acceptance [1], are implemented; (3) candidate acceptance, where validated candidates are integrated into the input and KV cache is updated accordingly. The inference scheme in Figure 2 illustrates the generation and verification combined in a single forward pass.
@@ 3.1 Prompt Tokens
The prompt tokens are the key component of PPD to realize multi-token generation. Initially introduced in [12] to adapt LLMs for specific tasks, prompt tokens are typically prepended to the input, with outputs generated in an autoregressive manner. In this work, we propose a novel approach of utilizing prompt tokens by strategically positioning them at locations where tokens are anticipated to be generated in parallel. For conventional parallel decoding techniques [23, 1] that presume complete conditional independence among tokens decoded in a single step, the exact conditional probability is approximated by , where is the input, represents the generated outputs so far, indicates the token distance.333The token distance is the number of tokens between the last accepted token and the predicted token. However, we observe that as increases, the gap between the actual probability and its approximation expands, primarily due to the absence of relevant conditional dependencies. We argue that prompt tokens can bridge this gap by more accurately modeling the conditional probability as where is the prompt token with token distance . Through this forward pass in the decoder layers, these causally linked prompt tokens facilitate the flow of information along the sequence of speculative tokens, thus restoring the conditional probability.
@@ 3.2 Ensemble Prompt Tokens
Inspired by prompt ensembling [12], which uses multiple prompts to generate diverse responses and aggregates these to derive a single answer, we introduce the concept of ensemble prompt token (EPT). This additional abstraction allows us to decouple each prompt token from the fixed embedding dimension. For every prompt token, there exist multiple corresponding EPTs, each with its distinct embedding. We modify the attention mask to ensure that each EPT only depends on the corresponding EPTs from preceding prompt tokens. This selective visibility is maintained for both training and inference, where the guess token for each prompt token is determined by averaging the logits of its EPTs. The use of EPTs not only enables direct and flexible control over the trainable parameters, but also leads to an increase in prediction accuracy. The probability is approximated as , where denotes the EPT at a token distance of . 444Further details about EPTs can be found in Appendix B.
@@ 3.3 Training
Knowledge Distillation: To align the predictive behavior of prompt tokens with the original LLM, we employ knowledge distillation. Instead of using hard labels, prompt tokens are trained against the logits produced by the original LLM. Following Medusa [1], The loss function is formulated as:
@@
4 Dynamic Sparse Tree
4.1 Motivation
To achieve higher speedup, PPD utilizes a specialized tree attention [1, 18] to process multiple candidates within a single decoding step without expanding the batch size. Notably, PPD employs a sparse tree [1, 4], designed to prioritize candidates with higher prediction accuracy. One key distinction from the sparse tree used in previous works is the appending of a sequence of prompt tokens to each tree node as shown in Figure 3. To optimize the amortized acceptance length across decoding steps, it is crucial to carefully balance the number of candidate tokens and prompt tokens. Instead of appending a uniform number of prompt tokens to every candidate token, we allocate them based on each candidate’s probability, causing the tree’s maximum depth and structure to vary at each decoding step.
4.2 Construction Algorithm
We aim to construct a dynamic sparse tree that maximizes the amortized number of tokens generated with limited candidate tokens and prompt tokens. We first define the tree construction algorithm as a constrained optimization problem.
Definition 4.1.
Let be the maximal number of prompt tokens per tree node. The dynamic sparse tree can exist in states, each represented by corresponding to state , where . Let denote the subtree of composed solely of candidate tokens. The maximum depth of is .
Proposition 4.1.
For a dynamic sparse tree state , where each candidate token follows a path from the root, and the acceptance probability at each path position , the expected number of tokens generated is given by , where represents the contribution of a token to the expected number of tokens.
Proposition 4.2.
The expected total number of tokens generated for the dynamic sparse tree state at the current and the next decoding step is given by , where represents the state transition probability from state to state .
We are now ready to introduce Proposition 4.3, which we use in the pruning algorithm.
Proposition 4.3.
For a dynamic sparse tree state with candidate subtree , the change in expected total tokens due to the removal of a prompt token at candidate token is given by , where is the acceptance probability of candidate , denotes the number of prompt tokens prior to removal. We assume that .
To construct an approximately optimal dynamic sparse tree with specified numbers of candidate and prompt tokens, the process includes: (1) Optimal Candidate Trees: Constructing trees using only candidate tokens at varying depths, employing the algorithm from Medusa [1] and Sequoia [4] to maximize as stated in Proposition 4.1; (2) Appending Prompt Tokens: Attaching the maximum allowable prompt tokens to each candidate token from the first step; (3) Greedy Prompt Token Removal: Removing prompt tokens greedily to minimize (Proposition 4.3), continuing until the desired prompt token count is reached.
Proposition 4.4.
The amortized number of tokens generated for the dynamic sparse tree state is given by , where is the steady-state probability of state , and is the function defined in Proposition 4.1.
Hardware-awareness. All the probabilities used above can be approximated on a validation dataset. The dynamic sparse tree construction algorithm can now be formulated as finding the dynamic sparse tree with candidate tokens and prompt tokens to maximize :

### v3

@@ Abstract
The auto-regressive decoding of Large Language Models (LLMs) results in significant overheads in their hardware performance. While recent research has explored various speculative decoding techniques for multi-token generation, these methods introduce high memory costs from the additional weights and KV cache of separate draft models, limiting efficiency in edge and long-context scenarios. To overcome these limitations in edge-scale LLMs, we propose a novel parallel prompt decoding that requires only runtime memory overhead by employing a unified single model for both speculation and verification. Inspired by the human natural language generation process, PPD approximates outputs generated at future timesteps in parallel by using multiple prompt tokens. Furthermore, we present a hardware-aware two-stage tree pruning algorithm that adaptively optimizes this decoding scheme to fully leverage the computational capacities on different GPUs. Through extensive experiments across LLMs ranging from MobileLlama to Vicuna-13B on a wide range of benchmarks, our approach demonstrates up to 2.49 speedup. Moreover, our parallel prompt decoding can serve as an orthogonal optimization for synergistic integration with existing speculative decoding, showing up to further speed improvement. To support future development, we have included our code implementation with this submission.
[…]
Hardware-Aware Parallel Prompt Decoding for
Memory-Efficient Acceleration of LLM Inference
††thanks: Accepted at EMNLP 2025
Hao (Mark) Chen1 Wayne Luk1 Ka Fai Cedric Yiu2 Rui Li3 Konstantin Mishchenko3 Stylianos I. Venieris3 Hongxiang Fan1,3 1Imperial College London, UK 2Hong Kong Polytechnic University, Hong Kong 3Samsung AI Center, Cambridge, UK {hc1620,w.luk}@ic.ac.uk {rui.li,s.venieris}@samsung.com konsta.mish@gmail.com cedric.yiu@polyu.edu.hk hongxiangfan@ieee.org
@@ 1 Introduction
The recent advances in large language models (LLMs) are increasingly shaping and influencing a wide range of AI applications. However, autoregressive generation, the de facto approach employed in LLM inference, suffers from inadequate hardware performance due to its inherent sequential nature (Stern et al., 2018). Speculative decoding (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2024), an emerging acceleration technique, employs a guess-and-verify framework for LLM inference, where a smaller draft model first predicts multiple tokens sequentially and then the original LLM verifies them in parallel. Despite its potential, speculative decoding is constrained by high training costs and memory overhead, as shown in Figure 1. Particularly, this paper reveals that the extra runtime memory required for draft model storage and the KV cache—which scales linearly with sequence length—limits its practicality in edge devices, mobile environments, and long-context scenarios. Therefore, given the growing need for user privacy and personalization, this paper focuses on accelerating edge-scale LLMs via a novel more memory- and cost-efficient acceleration solution.
Recent efforts have explored the possibility of generating multiple tokens in parallel without relying on a separate transformer draft model (Santilli et al., 2023). Approaches such as inserting additional decoding heads (Cai et al., 2024) and retrieving frequently used tokens (He et al., 2023) are employed to enhance performance. However, these methods either make strong conditional independence assumptions for tokens generated within a single step (Cai et al., 2024) or rely on unparameterized trained tokens that do not scale effectively (Lin et al., 2024). Therefore, they often suffer from low acceptance rates during inference.
@@
To alleviate the complexity and overhead associated with the use of draft models while maintaining a high acceptance rate, we propose Parallel Prompt Decoding (PPD), a novel architecture-agnostic and memory-efficient framework that adopts prompt tuning for non-autoregressive LLM inference. Inspired by the human natural language generation process where continuous words like common expressions and phrases are produced simultaneously, PPD introduces the use of ensemble prompt tokens, the meticulously trained embeddings, for multi-token prediction. Specifically, these trained tokens are appended to the original input sequence in parallel, enabling the concurrent generation of multiple output tokens in a single forward pass. The key intuition of PPD lies in the observation that if trained properly, prompt tokens appended to the input can approximate tokens generated at future timesteps, thereby partially recovering the missing conditional dependency information for multi-token generation. By strategically positioning trained prompt tokens, PPD achieves up to a 28% higher acceptance rate when predicting long-range tokens. Moreover, to facilitate the optimized implementation of PPD across different hardware platforms, we propose a hardware-aware two-stage tree pruning technique that adaptively refines the prompt structure during runtime based on the computational resources available on the specific hardware. As shown in Figure 1(a), PPD achieves a comparable speedup to the state-of-the-art speculative decoding approaches with negligible memory overhead and reduced training cost.
@@
To demonstrate the effectiveness of our approach, we evaluate PPD on MobileLLaMA (Chu et al., 2023), Vicuna-7b and Vicuna-13b (Chiang et al., 2023). Running on a single GPU using the A100-40GB and RTX 4090, our method achieves a speedup ratio for inference from 2.12 to 2.49 across a diverse range of popular datasets including MT-Bench, HumanEval, and GSM8K. Our experiments demonstrate that PPD not only achieves comparable throughput to the state-of-the-art speculative decoding method, but it also manages this with significantly fewer trainable parameters—specifically, 0.0002% of trainable parameters—and incurs only a minimal memory overhead (0.0004%), showcasing that PPD is remarkably cost- and memory-efficient. The training of prompt tokens can be completed in 16 hours using one A100 GPU, 8 hours using four GeForce RTX 3090 GPUs, compared to the 1-2 days on four A100 GPUs required for Eagle (Li et al., 2024b). Furthermore, since PPD does not require the modification of the original LLM or the addition of extra networks, it is highly adaptable and orthogonal to other decoding techniques. For instance, it can be effectively combined with a draft model to further reduce inference latency.
@@
A hardware-aware two-stage tree pruning technique that adaptively optimizes the prompt structure of PPD at runtime based on the available compute and memory resources, facilitating its efficient deployment on various hardware platforms.
@@ 2.1 Speculative Decoding
The guessing phase of speculative decoding adopts a lightweight draft model to generate multiple tokens at an increased speed (Kim et al., 2024). During the verification stage, the original LLM determines the acceptance of the guessed tokens. It is worth noting that both draft and original models still follow the auto-regressive inference scheme.
Building on the speculative decoding scheme, various studies have been conducted to further optimize its inference speed. To improve the accuracy of the draft model, Eagle (Li et al., 2024b) incorporates the hidden features into the draft model’s forward pass. Recently, Eagle-2 (Li et al., 2024a) enhances their approach using a context-aware dynamic tree construction. However, both Eagle and Eagle-2 utilize a separate draft model for multi-token generation, diverging fundamentally from our prompt decoding approach. Moreover, their dynamic tree construction scheme is an orthogonal technique to our two-stage tree pruning method. SpecInfer (Miao et al., 2024) adopts a tree-based speculative inference and verification scheme, improving the diversity of speculation candidates. Sequoia (Chen et al., 2024) optimizes the sparse tree structure of speculative decoding by considering the capability of the underlying hardware platforms. Our tree pruning algorithm differs from Sequoia by accounting for two types of tokens in the tree: prompt tokens and guess tokens, whereas Sequoia only considers guess tokens.
@@ 2.2 Parallel Decoding
To overcome the inherent limitations of autoregressive inference and the memory overhead associated with using a separate draft model, several attempts have been made to integrate both guessing and verification using one unified model. Medusa111We categorize Medusa as parallel decoding because it only adopts LM heads instead of separate models. (Cai et al., 2024) introduces language model (LM) heads at the final layer of the original LLM, facilitating the generation of multiple tokens in a single forward pass. It also utilizes tree attention masks in its verification process to increase speed even further. To enhance token drafting with retrieval-augmented generation (Karpukhin et al., 2020), Rest (He et al., 2023) introduce retrieval-based decoding tailored for specific scenarios. Inspired by Jacobi decoding (Santilli et al., 2023) that adopts multiple special tokens to accelerate machine translation, Lookahead Decoding (Fu et al., 2024) improves upon this method by generating parallel n-grams and employing a caching memory pool. To capture more information while using multiple special tokens at distinct positions, PaSS (Monea et al., 2023) and BiTA Lin et al. (2024) train additional token embeddings for parallel decoding. Compared with these parallel decoding methods, our proposed PPD introduces two key innovations that distinguish it: 1) PPD trains the embeddings of parameterized ensemble prompt tokens, 2) it utilizes a hardware-aware two-stage tree pruning algorithm for designing a sparse tree tailored to each hardware platform.
@@ 3 Parallel Prompt Decoding (PPD)
The primary advantage of PPD lies in training embeddings for prompt tokens rather than developing a separate model. Our method integrates three substeps into a single decoding step, following the guess-and-verify strategy: (1) candidate generation, where multiple candidate continuations222A candidate token, also referred to as a ”guess token”, is a draft token generated from a prompt token. are predicted by strategically inserting the prompt tokens into the input sequence. We adopt tree attention (Miao et al., 2024) to merge the processing of multiple candidates into a single forward pass; (2) candidate verification, where two verification schemes, exact matching (Fu et al., 2024) and typical acceptance (Cai et al., 2024), are implemented; (3) candidate acceptance, where validated candidates are integrated into the input and KV cache is updated accordingly. Figure 2 presents the inference scheme of combining generation and verification steps in a single forward pass.
@@ 3.1 Prompt Tokens
The prompt tokens are the key component of PPD to realize multi-token generation. Initially introduced by Lester et al. (2021) to adapt LLMs for specific tasks, prompt tokens are typically prepended to the input, with outputs generated in an autoregressive manner. In this work, we propose a novel approach of utilizing prompt tokens by strategically positioning them at locations where tokens are anticipated to be generated in parallel.
In the standard decoding process, the probability of predicting the next token is expressed as the conditional probability , where is the input prompt, are the tokens generated so far, and is the next token to be predicted. For conventional parallel decoding techniques (Stern et al., 2018; Cai et al., 2024) that presume complete conditional independence among tokens decoded in a single step, the exact conditional probability is approximated by , where indicates the token distance.333The token distance is the number of tokens between the last accepted token and the predicted token. However, we observe that as increases, the gap between the actual probability and its approximation expands due to the absence of relevant conditional dependencies. We argue that prompt tokens can bridge this gap by more accurately modeling the conditional probability as , where is the prompt token with token distance . Through this forward pass in the decoder layers, these causally linked prompt tokens facilitate the flow of information along the sequence of speculative tokens, restoring the conditional probability. We demonstrate the effectiveness of this approach in Section 5.2.
@@ 3.2 Ensemble Prompt Tokens
Inspired by prompt ensembling (Lester et al., 2021), which uses multiple prompts to generate diverse responses and aggregates these to derive a single answer, we introduce the concept of ensemble prompt token (EPT). This abstraction allows us to decouple each prompt token from the fixed embedding dimension. For every prompt token, there exist multiple corresponding EPTs, each with its distinct embedding. We modify the attention mask to ensure that each EPT only depends on the corresponding EPTs from preceding prompt tokens. This selective visibility is maintained for both training and inference, where the guess token for each prompt token is determined by averaging the logits of its EPTs. The use of EPTs not only enables direct and flexible control over the trainable parameters, but also leads to increased prediction accuracy. The probability is approximated as , where denotes the EPT at a token distance of . Further details can be found in Appendix F.
@@ 3.3 Training
Knowledge Distillation: To align the predictive behavior of prompt tokens with the original LLM, we employ knowledge distillation. Instead of using hard labels, prompt tokens are trained against the logits produced by the original LLM. Following Medusa (Cai et al., 2024), the loss function is formulated as:
@@
4 Sparse Tree Pruning
4.1 Customized Sparse Tree Attention
Tree attention, introduced by SpecInfer (Miao et al., 2024), increases the expected acceptance rate by considering the top-k candidates from a single decoding step. In their approach, the input is structured as a tree, where each level of the tree corresponds to a specific output position. An attention mask is applied to the tree-structured input, allowing the model to process multiple candidates efficiently without increasing the batch size.
To improve the efficiency and performance of LLM inference, this paper proposes a novel sparse tree customized for PPD, which prioritizes candidates in the tree structure with higher prediction accuracy. A key difference from previous works (Cai et al., 2024; Chen et al., 2024) is the appending of a sequence of prompt tokens to each guess token. The length of the prompt token sequence decides the maximum depth of the speculative tree at the next decoding step. To further hide the latency introduced by the extra prompt tokens, we propose a novel tree pruning algorithm (Section 4.2) that optimizes the number of prompt tokens.
4.2 Two-Stage Tree Pruning Algorithm
As depicted in Figure 3, our tree pruning algorithm consists of two stages: an offline static tree pruning phase and an online hardware-aware tree optimization phase. These two stages are applied subsequently to reduce the amount of computation involved in PPD multi-token generation.
Static Tree Pruning. The first stage, static tree pruning, is applied offline prior to runtime deployment. The goal is to reduce the number of prompt tokens in the tree to achieve the desired tree size. As shown on the left side of Figure 3, the tree pruning process consists of three key steps:
- 1.
2.
Prompt Tokens Appending : Attaching the maximum allowable prompt tokens to each candidate token from the first step.
3.
Greedy Prompt Token Removal: Removing a prompt token greedily to maximize expected amortized acceptance lengths, continuing until the desired prompt token budget is reached.
Each guess token in the tree is appended with a sequence of prompt tokens, with each prompt token corresponding to a unique output position. The length of this sequence determines the tree’s maximum depth at the next decoding step. Thus, removing a prompt token at a guess token reduces the maximum tree depth at the next decoding step if this guess token is accepted in the current step. Let represent the acceptance probability of guess token , and denote the expected acceptance length with prompt tokens before removal. The decrease in expected acceptance length, , due to removing a prompt token at is given by . More details are discussed in Appendix B.
Hardware-Awareness Tree Optimization. Given that hardware platforms differ in terms of memory and compute resources, we propose a hardware-aware tree optimization to maximize the performance of PPD. As shown on the right of Figure 3, this optimization adjusts the tree size budget based on the performance characteristics of the target hardware.